# An Informative Planning Framework for Target Tracking and Active Mapping in Dynamic Environments with ASVs

### v2

@@ Abstract
Mobile robot platforms are increasingly being used to automate information gathering tasks such as environmental monitoring. Efficient target tracking in dynamic environments is critical for applications such as search and rescue and pollutant cleanups. In this letter, we study active mapping of floating targets that drift due to environmental disturbances such as wind and currents. This is a challenging problem as it involves predicting both spatial and temporal variations in the map due to changing conditions. We propose an informative path planning framework to map an arbitrary number of moving targets with initially unknown positions in dynamic environments. A key component of our approach is a spatiotemporal prediction network that predicts target position distributions over time. We propose an adaptive planning objective for target tracking that leverages these predictions. Simulation experiments show that our proposed planning objective improves target tracking performance compared to existing methods that consider only entropy reduction as the planning objective. Finally, we validate our approach in field tests using an autonomous surface vehicle, showcasing its ability to track targets in real-world monitoring scenarios.
@@ I Introduction
Target search and monitoring with autonomous mobile robots has many applications, such as search and rescue [1, 2], cleaning pollutants [3, 4], and wildlife monitoring [5]. Mobile robot platforms such as autonomous surface vehicles (ASVs), and unmanned aerial vehicles (UAVs) allow for faster and more efficient search [6] compared to traditional approaches such as manned or teleoperated surveys [7] and static sensor networks [8]. For target tracking in a priori unknown dynamic environments, a robot must trade off between exploring previously unseen regions to discover new targets and acquiring new observations of previously observed moving targets to update their position estimates. Monitoring in such dynamic environments remains a challenging problem.
@@
In this work, we consider the tracking of freely drifting targets with an ASV. We aim to build and maintain an accurate map of target positions in a dynamic environment for any number of targets with unknown initial positions. Partially submerged objects are subject to external disturbances such as waves, wind, and currents in the marine environment. This is a pertinent issue for maritime search and rescue [2], cleanups of plastic waste [3], or other pollutants [4].
@@
Figure 1: Our informative path planning (IPP) framework applied to monitoring freely drifting targets (colored buoy markers) with an ASV that is equipped with a stereo camera. A dynamic mapping strategy is used for target tracking. We introduce a spatiotemporal prediction network and a new IPP objective that enable accurate target tracking.
@@
Methods for finding sampling paths that maximize the information gathered for resource-constrained robots are termed informative path planning (IPP) [9]. Informative trajectories are quantified by an information-theoretic measure representing the mission objective, such as entropy in the map. For static environments, entropy is generally used as the planning objective for efficient spatial coverage [1, 10]. In dynamic environments, predicting spatial and temporal variations in the map over a certain time horizon is essential to plan informative trajectories for the robot. However, stochastic environmental disturbances such as wind and currents are hard to predict accurately for longer time frames.
Predicting changes in the environment is necessary to maintain an accurate map that represents the belief of target positions. Classical target tracking methods involve detecting, identifying, and estimating the velocities of multiple targets [11]. Using occupancy grid mapping for target tracking [12, 13] provides a simpler alternative by handling occupancy probabilities of the targets in a global probabilistic map representation. However, prior work [12, 13, 14] on occupancy grid mapping in dynamic environments learns state transitions through a series of observations, which is a challenging problem. If the dynamics governing the targets’ motion are known, they can be used to predict these changes.
@@
To overcome these limitations, we present an IPP framework for actively mapping moving targets in dynamic environments. We introduce a spatiotemporal prediction network to predict future target position distributions, taking into account the uncertain drift due to stochastic environmental forces. This enables more efficient prediction of the uncertainty in target positions for longer time horizons and works for any number of targets. Furthermore, our proposed planning objective explicitly considers tracking of targets with uncertain position estimates using predictions from the network. A dynamic occupancy grid mapping approach is used to update the map with the sensor observations and an estimate of the targets’ drift using the measured wind velocity at each time step.
@@
We make the following claims supported by extensive simulation experiments and field trials of our IPP framework implemented on an ASV: (i) our spatiotemporal network and the proposed IPP objective improve target tracking performance when compared to existing methods that use only entropy reduction as the planning objective, and (ii) incorporating drift predictions in the mapping stage improves mapping accuracy when compared to mapping without predictions. We open-source our software at https://github.com/sanjeevrs2000/ipp_dyntrack.
@@ II Related Work
Target search is often formulated as an IPP problem [9]. Various studies have investigated IPP approaches for mapping discrete and static targets [1, 15]. Barrionuevo et al. [3] use reinforcement learning for planning to monitor and collect floating waste with a fleet of two ASVs, assuming perfect localization of the targets. Chen and Liu [16] study IPP for monitoring in dynamic environments with a model-predictive tree search with multi-objective optimization. They, however, assume that the environmental dynamics are deterministic, when in reality, long-term predictions are difficult due to the stochastic nature of real-life environments. Similar to [16], we use a sampling-based tree search algorithm for IPP that incorporates both the spatial and temporal variations in the map. In addition, our proposed spatiotemporal prediction network informs the informative planner of the uncertainties in target positions over the planning horizon, taking into account the stochastic nature of the wind and currents.
Some studies use high-fidelity numerical modeling of the underlying physical processes to guide adaptive sampling of oceanographic processes [17, 18, 19]. However, their adaptive planning strategies are limited to myopic behaviors, i.e. single-step lookahead, due to computational and operational constraints [17, 18]. Similar to these previous works, we use a learning-based spatiotemporal prediction model to inform the adaptive planning strategy. Our spatiotemporal network allows us to efficiently estimate the target position distributions over longer periods.
@@
The occupancy grid mapping algorithm [20] is widely used in robotics to create probabilistic occupancy maps for static environments. Several studies extend the occupancy grid formulation for mapping in dynamic environments, as relevant to our problem setup. Meyer et al. [14] propose a framework for mapping in dynamic environments where the state transition probabilities are learned. The Bayesian occupancy filter [12] introduces a dynamic occupancy grid mapping method for multi-object tracking. The map is updated in two steps, namely the estimation and prediction steps, similar to Bayesian state estimation. While our method is similar to that of Coue et al. [12], we infer the measured wind speed in the prediction at each time step rather than learning it through observations. It allows us to maintain a map that is agnostic to the number of targets while retaining the probabilistic formulation. This dynamic occupancy grid, combined with predictions from the spatiotemporal network, provides a framework to tackle the exploitation-exploration trade-off in dynamic environments.
@@
Various recent studies apply deep learning-based methods for mapping temporally varying environments. Recurrent neural networks have been used to predict future occupancy states in dynamic environments [21, 22] in autonomous driving scenarios. Such methods rely on continuous observations of the targets of interest and predict the future dynamic states in the map for only a few seconds. Casado et al. [4] use a variational autoencoder to predict spatiotemporal pollutant distributions. Wang et al. [23] employ an attention-based neural network for persistent monitoring of mobile targets. In contrast to these works, we propose a physics-informed spatiotemporal prediction model that is independent of the number of targets and efficiently estimates uncertain target position distributions, for longer horizons of up to $30$ s. Overall, our approach enables efficient tracking of moving targets with unknown initial positions in dynamic environments by providing an accurate map of target positions and predicting the changes in the map due to the dynamic environment. This enables effective autonomous monitoring for applications such as search and rescue and pollutant cleanups.
@@ III Problem Statement
Our aim is to detect and actively map targets of interest, such as floating waste, influenced by environmental disturbances in an a priori unknown environment. We frame this as an IPP problem [9]. We seek a trajectory $\mathcal{T}$ from the set of all feasible trajectories $\Psi$, that maximizes an information-theoretic measure or utility $\mathcal{I(\mathcal{T})}$, given a limited budget B{B}, such as mission time. In particular, the goal is to approximate the optimal trajectory $\mathcal{T}^{*}$ given by
@@
where $C(\mathcal{T})$ is the cost associated with the trajectory $\mathcal{T}$.
@@ IV-A Overview
We introduce a novel IPP framework for actively mapping moving targets in dynamic environments. An overview of our framework is shown in Fig. 2. We consider an ASV with a stereo camera for object detection and localization. A probabilistic occupancy grid is used to map these detections. We also incorporate a prediction step in the mapping to compensate for the targets’ drift. A key aspect of our approach is a spatiotemporal prediction network to estimate the spatial and temporal variations of the target positions over future time steps, given the current state of the occupancy grid and the current wind conditions. We then propose a new IPP utility formulation for target tracking that leverages predictions from our spatiotemporal network.
@@ IV-B Perception
The ASV is equipped with a forward-facing stereo camera to detect and localize targets of interest. In this work, we use YOLOv8 [24] to identify the targets of interest, noting that our approach is agnostic to the choice of object detection algorithm. The positions of detected targets are projected onto the global occupancy map:
@@
where $\boldsymbol{p}_{c}$ is the estimated depth of the target in the camera frame, $\boldsymbol{p}_{g}$ is its position in the global east-north-up (ENU) frame. The rotation matrix for transformation from the camera to the body frame is denoted by $R_{bc}$, assuming the camera is forward-facing and the distance between the body’s origin and the camera is negligible. The body to global frame rotation matrix is given by $R_{z\psi}$, where $c$ and $s$ denote $\cos\psi$ and $\sin\psi$. The estimated global position and heading of the ASV are denoted by $(x_{v},y_{v})$ and $\psi$ respectively.
@@
Occupancy grid mapping requires inverse sensor models to convert sensor observations to belief probabilities on the map. To define inverse sensor models for our application, we perform an empirical analysis in simulations to estimate the localization error of the detected objects. The inverse sensor models we use are similar to previous work [10, 25], but are kept asymmetric to better suit our problem setup in dynamic environments. Another distinction is that the inverse sensor model for detection also associates an occupancy probability for neighboring cells of each detection with a Gaussian kernel to take into account the localization errors in the detections. We derive the standard deviation $\sigma(r)$ of this error as a function of the distance to the detection $r$, through empirical analysis. To do this, we place a single object at various distances and relative heading angles to the camera in the simulation environment. We show the resulting standard deviations in the errors of the estimated positions of the targets in Fig. 3. We fit a curve over this plot to estimate the error as a function of the distance. Our inverse sensor models are defined as:
$p\!\left(m_{i,j}\mid z\right)$ $=\frac{1}{1+e^{a(r-d)}}\cdot e^{-\tfrac{\lVert\boldsymbol{x}_{c}-\boldsymbol{x}_{d}\rVert^{2}}{2\sigma^{2}}}$ (3)
@@
where the standard deviation $\sigma(r)$ is set to $0.0012r^{2}$, based on the analysis in Fig. 3. The estimated position of the detected target is given by $\boldsymbol{x}_{d}$, and $\boldsymbol{x}_{c}-\boldsymbol{x}_{d}||$ denotes the distance of a nearby cell $c$ to this detection. The inverse sensor model associating a detection probability to cell $(i,j)$ given a detection is denoted by $p(m_{i,j}|z)$, parameterized by $a,d$. The negative inverse sensor model for updating free cells is given by $p(m_{i,j}|z^{\prime})$, parameterized by $a^{\prime},d^{\prime}$. The radial field of view is clipped at $35$ m, as depth measurements become unreliable for further detections. The inverse observation models are depicted in Fig. 3.
@@ IV-C Dynamic Occupancy Grid Mapping
We introduce a dynamic occupancy grid method for mapping, which is updated in two stages at each time step, namely the estimation and prediction steps. First, in the estimation step, the map is updated with the sensor observations, using the classical probabilistic occupancy grid algorithm [20]. In the prediction step, we compensate for the drift of targets in the map with the instantaneous wind speed. In the estimation step, we update the map recursively as:
@@
where $z_{t}$ is the sensor observation at time step $t$. The log odds nation of the occupancy state $p\left(m_{i,j}\mid z_{1:t}\right)$ of a cell $(i,j)$ given the observations $z_{1:t}$ up to $t$, is denoted by $l(m_{i,j}|z_{1:t})$. The log odds representation of the inverse sensor model $p(m_{i,j}|z_{t})$ is denoted by $l(m_{i,j}|z_{t})$.
In the prediction step, the occupancy grid is updated each time step to account for the drift of the targets. We consider targets of interest such as floating waste on the water surface. For such objects, the wave effects are ignored as the drift due to the wind is more significant [26, 27]. We assume that the drift velocity can be approximated as being proportional to the instantaneous wind velocity, in line with previous work [26]. The drift for a single time step can therefore be approximated as:
@@
where $v_{w}$ is the wind speed, $\psi_{w}$ is the direction of the wind velocity at time $t$, and $\Delta t$ is the time step between successive updates of the map. The wind factor $\gamma$ is defined as the ratio of the drift speed of the object to the wind speed, which we set to be $0.03$ following Rohrs et al. [26]. To update the occupancy grid with the drift prediction of the targets, we first filter it in order to keep only occupied cells (>0.5>0.5) as:
@@
where $m^{\prime}_{i,j}$ denotes the occupancy probabilities in the filtered grid $\mathcal{M^{\prime}}$. As the incremental drift at each time step is much smaller than the resolution of the occupancy grid, we store the accumulated drift in variables $R_{x},R_{y}$. In the prediction step, the occupancy probabilities are updated only if $\texttt{round}(R_{x})>0$ or $\texttt{round}(R_{y})>0$. The update rule for the prediction step is described in lines 20-22 of Algorithm 1, where $\alpha$ and $\beta$ are parameters to trade-off between the current occupancy state and the prediction. The occupancy probabilities in the grid are limited between $(p_{\mathrm{low}},p_{\mathrm{high}})=(0.15,\,0.9)$ to adapt the mapping to the dynamic setting. A condensed form of this dynamic occupancy grid mapping algorithm is given in Algorithm 1.

### v3

@@ Abstract
Mobile robot platforms are increasingly being used to automate information gathering tasks such as environmental monitoring. Efficient target tracking in dynamic environments is critical for applications such as search and rescue and pollutant cleanups. In this letter, we study active mapping of floating targets that drift due to environmental disturbances such as wind and currents. This is a challenging problem as it involves predicting both spatial and temporal variations in the map due to changing conditions. We introduce an integrated framework combining dynamic occupancy grid mapping and an informative planning approach to actively map and track freely drifting targets with an autonomous surface vehicle. A key component of our adaptive planning approach is a spatiotemporal prediction network that predicts target position distributions over time. We further propose a planning objective for target tracking that leverages these predictions. Simulation experiments show that this planning objective improves target tracking performance compared to existing methods that consider only entropy reduction as the planning objective. Finally, we validate our approach in field tests, showcasing its ability to track targets in real-world monitoring scenarios.
@@ I Introduction
Target search and monitoring with autonomous mobile robots has many applications, such as search and rescue [meera2019obstacle, ai2019intelligent], cleaning pollutants [barrionuevo2025optimizing], and wildlife monitoring [tokekar2013tracking]. Mobile robot platforms such as autonomous surface vehicles (ASVs), and uncrewed aerial vehicles (UAVs) allow for faster and more efficient search [dunbabin2012robots] compared to traditional approaches such as manned or teleoperated surveys [murphy2017disaster] and static sensor networks [lanzolla2021wireless]. For target tracking in a priori unknown dynamic environments, a robot must trade off between exploring previously unseen regions to discover new targets and acquiring new observations of previously observed moving targets to update their position estimates. Monitoring in such dynamic environments remains a challenging problem.
@@
In this work, we consider the tracking of freely drifting targets with an ASV. We aim to build and maintain an accurate map of target positions in a dynamic environment for any number of targets with unknown initial positions. Partially submerged objects are subject to external disturbances such as wind and currents in the marine environment. This is a pertinent issue for maritime search and rescue [ai2019intelligent], cleanups of plastic waste [barrionuevo2025optimizing], or other pollutants [wang2019dynamic].
@@
Figure 1: Our informative path planning (IPP) framework applied to monitoring freely drifting targets (colored buoy markers) with an autonomous surface vehicle that is equipped with a stereo camera. A dynamic mapping strategy is used for target tracking. We introduce a spatiotemporal prediction network and a new IPP objective that enable accurate target tracking.
@@
Methods for finding sampling paths that maximize the information gathered for resource-constrained robots are termed informative path planning (IPP) [popovic2024learning]. Informative trajectories are quantified by an information-theoretic measure representing the mission objective, such as entropy in the map. For static environments, entropy is generally used as the planning objective for efficient spatial coverage [meera2019obstacle, popovic2020informative]. In dynamic environments, predicting the spatial and temporal variations in the map and frequent replanning are required to maintain an accurate map of target positions.
Some studies use numerical models of known physical processes to map the environment [wang2019dynamic, du2021multi]. In contrast, we focus on tracking discrete targets that drift, which are difficult to model due to the stochastic nature of environmental disturbances [breivik2011wind]. Classical target tracking methods detect, associate identifiers with targets, and track them individually [robin2016multi]. Thus, the computational demand also scales with the number of targets. Using occupancy grid mapping for target tracking [coue2006bayesian] provides a simpler alternative by handling occupancy probabilities of the targets in a global probabilistic map representation, avoiding maintaining individual tracks and associating identifiers to targets. Prior works [wolf2004online, meyer2012occupancy, coue2006bayesian] on occupancy grid mapping in dynamic environments learn state transitions online, but such methods can become unsuitable for real-time applications due to computational constraints.
@@
To overcome these limitations, we present an integrated solution for actively mapping moving targets in dynamic environments. A dynamic occupancy grid mapping approach is used to update the map, combining the sensor observations and an approximation of the targets’ drift using the measured wind velocity at each time step. We employ adaptive sampling-based planning to enable the robot to track targets by driving it towards regions of interest. We introduce a spatiotemporal prediction network to estimate uncertain target positions for any number of targets. This enables efficient predictions for longer time horizons, based only on the current map state and the measured wind speed. Furthermore, our proposed planning objective explicitly considers target tracking uncertainty by using target position predictions from the network.
@@
We make the following claims supported by extensive simulation experiments and field trials of our IPP framework implemented on an ASV: (i) Our proposed IPP objective that utilizes the spatiotemporal network improves target tracking performance when compared to existing methods that use only entropy reduction as the planning objective, and (ii) incorporating drift predictions in the mapping stage improves mapping accuracy when compared to mapping without predictions. As a result, our framework enables efficient real-time adaptive planning and active mapping in dynamic environments, as also validated through field trials. We open-source our software at https://github.com/sanjeevrs2000/ipp_dyntrack.
@@ II Related Work
Various studies have investigated IPP approaches for mapping discrete and static targets [meera2019obstacle, vashisth2024deep, moon2022tigris]. Barrionuevo et al. [barrionuevo2025optimizing] use reinforcement learning for planning to monitor and collect floating waste with a fleet of two ASVs, assuming perfect localization of the targets. Chen and Liu [chen2019multi] study IPP for monitoring in dynamic environments with a model-predictive tree search with multi-objective optimization. They, however, assume the environmental dynamics to be deterministic when in reality, long-term predictions are difficult due to the stochastic nature of real-life environments. Similar to [chen2019multi], we use a sampling-based planning approach for IPP that incorporates both the spatial and temporal variations in the map in the planning objective. In addition, our proposed spatiotemporal prediction network informs the informative planner of the uncertainties in target positions over the planning horizon, taking into account the stochastic nature of the wind and currents.
Some literature in robot monitoring for source seeking and plume tracking uses high-fidelity numerical models of the underlying physical processes [preston2024phortex, wang2019dynamic, du2021multi, ma2025adaptive]. However, their adaptive planning strategies are often limited to myopic behaviors, i.e., single-step lookahead, due to computational and operational constraints. A body of existing work also addresses robot monitoring in dynamic environments with multiple robots [du2021multi, wang2019dynamic]. Dividing roles between a team of robots simplifies decision making when there are multiple objectives [wang2019dynamic, barrionuevo2025optimizing]. In contrast, we focus on tracking drifting targets with a single ASV while noting that most of our framework could still be applicable in a multirobot setting.
@@
The occupancy grid mapping algorithm [elfes1989using] is widely used in robotics to create probabilistic occupancy maps for static environments. Several studies extend the occupancy grid formulation for mapping in dynamic environments, as relevant to our problem setup. Meyer et al. [meyer2012occupancy] propose a framework for mapping in dynamic environments where the state transition probabilities are learned. Wolf and Sukhatme [wolf2004online] study mapping in dynamic environments by maintaining separate maps for static and dynamic parts. The Bayesian occupancy filter [coue2006bayesian] introduces a dynamic occupancy grid mapping method for multi-object tracking. The map is updated in two steps, namely the estimation and prediction steps, similar to Bayesian state estimation. While our method is similar to that of Coue et al. [coue2006bayesian], we infer the measured wind speed for prediction at each time step rather than learning it through observations. It allows us to maintain a map that is agnostic to the number of targets while retaining the probabilistic formulation. This dynamic occupancy grid, combined with predictions from the spatiotemporal network for efficient replanning, provides a framework capable of real-time mapping and planning.
@@
Various recent studies apply deep learning-based methods for mapping temporally varying environments. Recurrent neural networks have been used to predict future occupancy states in dynamic environments in autonomous driving scenarios [mann2022predicting, toyungyernsub2021double]. Such methods rely on continuous observations of the targets of interest and predict the future dynamic states in the map for only a few seconds. Wang et al. [wang2023spatio] employ an attention-based neural network for persistent monitoring of mobile targets. In contrast to these works, we propose a physics-informed spatiotemporal prediction model that is independent of the number of targets and efficiently estimates uncertain target position distributions, for longer horizons of up to $30$ s. Overall, our approach enables efficient tracking of moving targets with unknown initial positions in dynamic environments. This enables effective autonomous monitoring for applications such as search and rescue and pollutant cleanups.
@@ III Problem Statement
Our aim is to detect and actively map targets of interest, such as floating waste, influenced by environmental disturbances in an a priori unknown environment. We frame this as an IPP problem [popovic2024learning]. We seek a trajectory $\mathcal{T}$ from the set of all feasible trajectories $\Psi$, that maximizes an information-theoretic measure or utility $I(\mathcal{T})$, given a limited budget B{B}, such as mission time. In particular, the goal is to approximate the optimal trajectory $\mathcal{T}^{*}$ given by
@@
where $C(\mathcal{T})$ is the cost associated with the trajectory $\mathcal{T}$. In this work, we consider constant speed trajectories. The trajectory planning problem thus reduces to searching for a geometric path that maximizes $I(\mathcal{T})$.
@@ IV-A Overview
We introduce a novel IPP framework for actively mapping moving targets in dynamic environments. An overview of this framework is shown in Fig. 2. We consider an ASV with a stereo camera for object detection and localization. A probabilistic occupancy grid is used to map these detections. We also incorporate a prediction step in the mapping to compensate for the targets’ drift. A key aspect of our approach is a spatiotemporal prediction network to estimate the spatial and temporal variations of the target positions over future time steps, given the current state of the occupancy grid and the current wind conditions. We then propose a new IPP utility formulation for target tracking that leverages predictions from our spatiotemporal network.
@@ IV-B Perception
The ASV is equipped with a forward-facing stereo camera to detect and localize targets of interest. In this work, we use YOLOv8 [yolov8_ultralytics] to identify the targets of interest, noting that our approach is agnostic to the choice of object detection algorithm. The positions of detected targets are projected onto the global occupancy map:
@@
where $\boldsymbol{p}_{c}$ is the estimated depth of the target in the camera frame, $\boldsymbol{p}_{g}$ is its position in the global east-north-up (ENU) frame. The rotation matrix for transformation from the camera to the body frame is denoted by $R_{bc}$, assuming the camera is forward-facing and the distance between the body’s origin and the camera is negligible. The body to global frame rotation matrix is given by $R_{z\psi}$, where $c$ and $s$ denote $\cos\psi$ and $\sin\psi$. The estimated global position and heading of the ASV are denoted by $(x_{v},y_{v})$ and $\psi$, respectively.
@@
Occupancy grid mapping requires inverse sensor models to convert sensor observations to belief probabilities on the map. To define inverse sensor models for our application, we perform an empirical analysis in simulations to estimate the localization error of the detected objects. The inverse sensor models we use are similar to previous work [popovic2020informative, moon2022tigris], but are kept asymmetric to better suit our problem setup in dynamic environments. Another distinction is that the inverse sensor model for detection also associates an occupancy probability for neighboring cells of each detection with a Gaussian kernel to take into account the localization errors in the detections. We derive the standard deviation $\sigma(r)$ of this error as a function of the distance to the detection $r$, through empirical analysis. To do this, we place a single object at various distances and relative heading angles to the camera in the simulation environment. We show the resulting standard deviations in the errors of the estimated positions of the targets in Fig. 3. A curve is fit over this plot to estimate error as a function of the distance. Our inverse sensor models, for detections ($z$), and free cells ($z^{\prime}$), are defined as:
$p\!\left(m_{i,j}\mid z\right)$ $=\frac{1}{1+e^{a(r-d)}}\cdot e^{-\tfrac{\lVert\boldsymbol{x}_{c}-\boldsymbol{x}_{d}\rVert^{2}}{2\sigma^{2}}},$ (3)
@@
where the standard deviation $\sigma(r)$ is set to $0.0012r^{2}$, based on the analysis in Fig. 3. The estimated position of the detected target is given by $\boldsymbol{x}_{d}$, and $\boldsymbol{x}_{c}-\boldsymbol{x}_{d}||$ denotes the distance of a nearby cell $c$ to this detection. The inverse sensor model associating a detection probability to cell $(i,j)$ given a detection is denoted by $p(m_{i,j}|z)$, parameterized by $a,d$. The negative inverse sensor model for updating free cells is given by $p(m_{i,j}|z^{\prime})$, parameterized by $a^{\prime},d^{\prime}$. The radial field of view (FOV) is clipped at $35$ m, as depth measurements become unreliable for further detections. The inverse observation models are depicted in Fig. 3.
@@ IV-C Dynamic Occupancy Grid Mapping
We introduce a dynamic occupancy grid method for mapping to keep track of detected target positions. The map is updated in two stages at each time step, namely the estimation and prediction steps. First, in the estimation step, the map is updated with the sensor observations, using the classical occupancy grid algorithm [elfes1989using]. In the prediction step, we compensate for the drift of targets in the map with the instantaneous wind speed. In the estimation step, we update the map recursively as:
@@
where $z_{t}$ is the sensor observation at time step $t$. The log odds notation, $l(.)=\log\big(p(.)/(1-p(.))\big)$, of the occupancy probability of cell $(i,j)$, $p(m_{i,j})$, is denoted by $l(m_{i,j})$. The log odds representation of the inverse sensor model $p(m_{i,j}|z_{t})$ is denoted by $l(m_{i,j}|z_{t})$.
In the prediction step, the occupancy grid is updated each time step to account for the drift of the targets. We assume that the drift velocity of small floating objects, such as the ones we consider, can be approximated to be proportional to the instantaneous wind velocity, in line with previous work [rohrs2015drift]. The drift for a single time step can therefore be approximated as:
@@
where $v_{w}$ is the wind speed, $\psi_{w}$ is the direction of the wind velocity at time $t$, and $dt$ is the time step between successive updates of the map. The wind factor $\gamma$ is defined as the ratio of the drift speed of the object to the wind speed, which we set to be $0.03$ following [rohrs2015drift]. To update the occupancy grid with the drift prediction of the targets, we first filter it in order to keep only occupied cells (>0.5>0.5) as:
@@
where $m^{\prime}_{i,j}$ denotes the occupancy probabilities in the filtered grid $\mathcal{M^{\prime}}$. As the incremental drift at each time step is much smaller than the resolution of the occupancy grid, we store the accumulated drift in variables $R_{x},R_{y}$. In the prediction step, the occupancy probabilities are updated only if $\texttt{round}(R_{x})>0$ or $\texttt{round}(R_{y})>0$. The update rule for the prediction step is described in lines 20-22 of Algorithm 1, where $\alpha$ and $\beta$ are parameters to trade off between the current occupancy state and the prediction. The occupancy probabilities in the grid are constrained to the interval $p_{\mathrm{low}},p_{\mathrm{high}}]=[0.15,\,0.9]$ to adapt the mapping to the dynamic setting. A condensed form of this dynamic occupancy grid mapping algorithm is given in Algorithm 1.